# OT Score: An OT based Confidence Score for Unsupervised Domain Adaptation

### v1

@@ Abstract
We address the computational and theoretical limitations of existing distributional alignment methods for unsupervised domain adaptation (UDA), particularly regarding the estimation of classification performance and confidence without target labels. Current theoretical frameworks for these methods often yield computationally intractable quantities and fail to adequately reflect the properties of the alignment algorithms employed. To overcome these challenges, we introduce the Optimal Transport (OT) score, a confidence metric derived from a novel theoretical analysis that exploits the flexibility of decision boundaries induced by Semi-Discrete Optimal Transport alignment. The proposed OT score is intuitively interpretable, theoretically rigorous, and computationally efficient. It provides principled uncertainty estimates for any given set of target pseudo-labels without requiring model retraining, and can flexibly adapt to varying degrees of available source information. Experimental results on standard UDA benchmarks demonstrate that classification accuracy consistently improves by identifying and removing low-confidence predictions, and that OT score significantly outperforms existing confidence metrics across diverse adaptation scenarios.
@@ 1 Introduction
In recent years, deep neural networks have achieved remarkable breakthroughs across a wide range of applications. However, if the distribution of the training and test data differs, significant performance degradation occurs, which is known as a domain shift [29], which makes retraining critical for the model to re-gain the generalization ability in new domains.
Unsupervised domain adaptation (UDA) mitigates the domain shift problem where only unlabeled data is accessible in the target domain [9]. A key approach for UDA is aligning the distributions of both domains by mapping data to a shared latent feature space. Consequently, a classifier trained on source domain features in this space can generalize well to the target domain. Several existing works [16], [17], [4], [3], [25] exhibits a principled way to transform target distribution to be ’closer’ to the source distribution so that the classifier learned from the source data can be directly applied to the target domain thus pseudo-labels (or predictions) can be made accordingly.
This leads to the question of whether such transformations from the target to the source distribution can accurately match the corresponding class-conditional distributions. For any given target dataset, it is always possible to align its feature distribution with that of the source domain using a divergence function, regardless of whether classes overlap. However, performing UDA in this way is reasonable only if target features remain well-separated by the decision boundaries induced through alignment in the latent feature space—something that is typically difficult to determine in practice. Moreover, the distributional alignment approach complicates the identification of samples with low-confidence pseudo-labels (i.e., samples close to overlapping regions), potentially causing noisy supervision and thus degrading classification performance. This issue becomes particularly critical when no labeled information for the target data is available. To address these challenges, we will focus on the following two questions in this work:
Question 1: What kind of domain shift can be aligned back with the source distribution while preserving the correct class labels?
Question 2: Is there a theoretically guaranteed confidence score that can be used to evaluate the uncertainty of given pseudo-labels?
To answer these questions, it is necessary to specify which distance is used for the distributional alignment since the pulling strategy on the target varies according to the divergence used. Optimal Transport (OT) distance is a popular distance adopted in UDA problems [4], [3], [25]. Using OT distance has an intuitive interpretation since it considers the overall geometry of data. Redko et al. [23] provides a theoretical guarantee connecting generalization errors on source and target domains based on the Wasserstein distance, which validates the idea of minimizing Wasserstein distance between source and target distributions:
Theorem 1 (Informal [23]).
@@
Despite its advantages, the theoretical understanding of label-preserving alignment under OT loss remains underexplored. A bad pulling strategy on $\mathcal{T}$ might minimize $W_{1}$ term to 00 without any guarantee for the $\lambda$ term in the feature space. We address this theoretical gap in Section 4.1, thereby answering Question 1.
@@
Guided by our theoretical analysis under the semi-discrete OT framework (Section 4.2), we address Question 2 by introducing the OT score—a confidence metric designed to quantify uncertainty in pseudo-labeled target samples. As illustrated in Figure 1, the OT score reflects the flexibility of decision boundaries induced by semi-discrete OT alignment, which enables effective uncertainty estimation in the target domain. This allows the algorithm to abstain from classifying samples with high uncertainty. Compared to fully continuous or fully discrete OT formulations, semi-discrete OT is computationally more efficient, especially in high-dimensional spaces and large-scale datasets. It provides a flexible framework that can be readily adapted to scenarios with limited source information (e.g., cases where only class-wise mean features from the source are available). Such scenarios are practical and realistic in real-world applications, as they facilitate the use of source information while preserving privacy and minimizing computational and storage overhead. Furthermore, our method is compatible with pseudo labels generated by state-of-the-art DA methods, which may also leverage partial source-domain information. A detailed comparison with existing confidence scores is provided in Section 2.
@@
[…]
2 Related Works
Theory about DA: Several theoretical works have investigated the learnability and generalization guarantees of domain adaptation (DA). Specifically, [1] analyzes the DA learnability problem and sample complexity under the standard VC-dimension framework, and identifies a setting in which no algorithm can successfully solve the DA problem. In a related direction, [24] provides a theoretical analysis about the existence of a hypothesis that performs well across both source and target domains, and further establishes finite-sample approximation properties of the $\lambda$ term. A closely related theoretical work is [14], which alleviates the label mismatching problem by searching for a transformation $T$ that satisfies the following conditions: (1) $T_{\#}\mu_{S}=\mu_{T}$, and (2) $T$ preserves the labels.
Confidence Scores: Uncertainty estimation and confidence score have been prevalently employed in machine learning to improve model robustness. In particular, ordinal ranking techniques have been commonly used for selective classification ([13]; [6]; [19]; [20]), where the goal is to prioritize or filter samples based on their confidence scores in order to exclude low-confidence samples during training. Lee et al. [15] propose the JMDS score to effectively identify low-confidence samples, thereby enhancing the reliability of the DA process. However, most existing confidence scores rely primarily on cluster-level information in the feature space, without explicitly modeling the geometric relationship between domains. In contrast, our proposed OT score take into account the geometry induced by the OT map, establishing a stronger connection between the source and target domains when computing confidence scores.
@@ 3.1 Domain Adaptation
Let $\Omega\subseteq\mathbb{R}^{d}$ be the sample space and $\mathcal{P}(\Omega)$ be the set of all probability measures over $\Omega$. In a general supervised learning paradigm for classification problems, we have a labeling function $f_{\bm{\theta}^{*}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ obtained from a parametric family $f_{\bm{\theta}}$ by training on a set of points $\mathbf{X}^{\mathcal{S}}=\{x_{1}^{\mathcal{S}},...,x_{N^{S}}^{\mathcal{S}}\}$ sampled from a source distribution $\mathbf{P}_{\mathcal{S}}\in\mathcal{P}(\Omega)$ and corresponding one-hot encoded labels $\mathbf{Y}^{\mathcal{S}}=\{y_{1}^{\mathcal{S}},...,y_{N^{\mathcal{S}}}^{\mathcal{S}}\}$.
Let $\mathbf{X}^{\mathcal{T}}=\{x_{1}^{\mathcal{T}},...,x_{N^{\mathcal{T}}}^{\mathcal{T}}\}$ be a dataset sampled from a target distribution $\mathbf{P}_{\mathcal{T}}\in\mathcal{P}(\Omega)$ without label information. The difference between $\mathbf{P}_{\mathcal{S}}$ and $\mathbf{P}_{\mathcal{T}}$ may lead to a poor performance if we use $f_{\bm{\theta}^{*}}$ for the new classification problem. In order to overcome the challenge of distributional shift, a common way is to decompose a neural network $f_{\bm{\theta}}$ into a feature mapping $\phi_{\bm{v}}$ composed with a classifier $h_{\bm{w}}$ such that $f_{\bm{\theta}}=h_{\bm{w}}\circ\phi_{\bm{v}}$, followed by minimizing the distance between $(\phi_{\bm{v}^{*}})_{\#}\mathbf{P}_{\mathcal{S}}$ and $(\phi_{\bm{v}})_{\#}\mathbf{P}_{\mathcal{T}}$ so that the target distribution will be aligned with the source distribution in the feature space. Then we may classify (or assign pseudo labels to) target data points based on the optimization result in the feature space. Various choices of divergence objective $D((\phi_{\bm{v}^{*}})_{\#}\mathbf{P}_{\mathcal{S}},(\phi_{\bm{v}})_{\#}\mathbf{P}_{\mathcal{T}})$ can be utilized. In this work, we focus on the distributional alignment between $(\phi_{\bm{v}^{*}})_{\#}\mathbf{P}_{\mathcal{S}}$ and $(\phi_{\bm{v}})_{\#}\mathbf{P}_{\mathcal{T}}$ using Wasserstein distance.
@@ 3.2.1 General Theory of OT
[…]
@@
[…]
is referred as the Kantorovitch problem in optimal transport. It is shown by Kantorovich–Rubinstein Duality theorem that (KP) has a dual form [26]:
@@
With the dual problem introduced, [2] proves Brenier’s theorem, which gives a sufficient condition under which the minimizer of the optimal transport problem is unique and is induced by a map $T=\nabla\phi$ for some convex function $\phi$, i.e. the OT map exists.
Under mild conditions on $\mu$ and $\nu$, Brenier’s theorem is satisfied when $c(x,y)=\|x-y\|_{p}$ for $p>1$. Although there is no guarantee about uniqueness of the optimal transport map when $p=1$, the existence of an optimal transport map can be proved through a secondary variational problem [26]:
@@
[…]
@@ 3.2.2 Semi-discrete Optimal Transport
[…]
@@ 4 Theoretical Analysis
In this section, we present theoretical insights into the use of OT for addressing DA problems. Complete proofs of all theoretical results are provided in Appendix B. For clarity and tractability, we focus on binary classification tasks. As discussed in Section 3.1, our interest lies in neural network–based DA. To this end, we adopt assumptions inspired by Neural Collapse [12], a prevalent phenomenon observed in well-trained neural networks. The extent to which the target feature distribution conforms to the Neural Collapse structure depends on the severity of the distributional shift between the source and target domains.
@@
Neural collapse (NC) is a phenomenon observed in well-trained neural networks where the learned features of samples belonging to the same class converge to a single point or form tightly clustered structures in the feature space, while the features of different classes become maximally separated. NC emerges while training modern classification DNNs past zero error to further minimize the loss [21]. During NC, the class means of the DNN’s last-layer features form a symmetric structure with maximal separation angle, while the features of each individual sample collapse to their class means. This simple structure of the feature layer not only appears beneficial for generalization but also helps in transfer learning and adversarial robustness. There are three main theoretical frameworks proposed to explain the emergence of NC: "Unconstrained Features Model" [18], [28], [11], "Local Elasticity" [31] and "Neural (Tangent Kernel) Collapse" [27].
In the following subsection, we focus on the setting where the class-conditional distributions in both the source and target domains are supported on, or concentrated within, bounded subsets of the feature space. Under this assumption, we analyze how data clusters are transported by the OT map.
@@ 4.1 Sufficient Conditions for Correct Classification
4.2 Post-check and OT Score
Table 1: Evaluation of confidence scores based on AURC (DSAN).
Dataset Task Maxprob Ent Cossim JMDS OT Score
Office-Home Ar $\rightarrow$ Cl 0.4306 0.4284 0.4170 0.4515 0.3403
Ar $\rightarrow$ Pr 0.2745 0.2738 0.2512 0.2849 0.2133
Ar $\rightarrow$ Rw 0.1469 0.1493 0.1521 0.1860 0.1157
Cl $\rightarrow$ Ar 0.2600 0.2631 0.2340 0.3228 0.2097
Cl $\rightarrow$ Pr 0.1757 0.1777 0.1612 0.2225 0.1503
Cl $\rightarrow$ Rw 0.1834 0.1848 0.1865 0.2246 0.1493
Pr $\rightarrow$ Ar 0.2371 0.2381 0.2245 0.2776 0.1984
Pr $\rightarrow$ Cl 0.3139 0.3105 0.3149 0.3302 0.2711
Pr $\rightarrow$ Rw 0.0974 0.0992 0.1037 0.1250 0.0817
Rw $\rightarrow$ Ar 0.1301 0.1318 0.1268 0.1751 0.1023
Rw $\rightarrow$ Cl 0.2581 0.2555 0.2641 0.2718 0.2112
Rw $\rightarrow$ Pr 0.0681 0.0684 0.0628 0.1026 0.0561
Avg. 0.2146 0.2150 0.2082 0.2478 0.1749
VisDA-2017 T $\rightarrow$ V 0.2301 0.2290 0.2289 0.2296 0.1799

### v2

# OT Score: An OT based Confidence Score for Source Free Unsupervised Domain Adaptation

@@ Abstract
We address the computational and theoretical limitations of current distributional alignment methods for source-free unsupervised domain adaptation (SFUDA). In particular, we focus on estimating classification performance and confidence in the absence of target labels. Current theoretical frameworks for these methods often yield computationally intractable quantities and fail to adequately reflect the properties of the alignment algorithms employed. To overcome these challenges, we introduce the Optimal Transport (OT) score, a confidence metric derived from a novel theoretical analysis that exploits the flexibility of decision boundaries induced by Semi-Discrete Optimal Transport alignment. The proposed OT score is intuitively interpretable and theoretically rigorous. It provides principled uncertainty estimates for any given set of target pseudo-labels. Experimental results demonstrate that OT score outperforms existing confidence scores. Moreover, it improves SFUDA performance through training-time reweighting and provides a reliable, label-free proxy for model performance.
@@ 1 Introduction
In recent years, deep neural networks have achieved remarkable breakthroughs across a wide range of applications. However, if the distribution of the training and test data differs, significant performance degradation occurs, which is known as a domain shift (Tsymbal, 2004), which makes retraining critical for the model to re-gain the generalization ability in new domains.
Unsupervised domain adaptation (UDA) mitigates the domain shift problem where only unlabeled data is accessible in the target domain (Glorot et al., 2011). A key approach for UDA is aligning the distributions of both domains by mapping data to a shared latent feature space. Consequently, a classifier trained on source domain features in this space can generalize well to the target domain. Several existing works (Long et al., 2015; 2017; Damodaran et al., 2018; Courty et al., 2016; Rostami & Galstyan, 2023) exhibit a principled way to transform target distribution to be "closer" to the source distribution so that the classifier learned from the source data can be directly applied to the target domain thus pseudo-labels (or predictions) can be made accordingly.
This leads to the question of whether such transformations from the target to the source distribution can accurately match the corresponding class-conditional distributions. For any given target dataset, it is always possible to align its feature distribution with that of the source domain using a divergence function, regardless of whether classes overlap. However, performing UDA in this way is reasonable only if target features remain well-separated by the decision boundaries induced through alignment in the latent feature space—something that is typically difficult to determine in practice. Moreover, the marginal distribution alignment approach complicates the identification of samples with low-confidence pseudo-labels (i.e., samples close to overlapping regions), potentially causing noisy supervision and thus degrading classification performance. This issue becomes particularly critical when no labeled information for the target data is available. Some existing works (Luo & Ren, 2021; Ge et al., 2023; Le et al., 2021) minimize a class-conditional discrepancy between the class-conditional feature distributions $P_{\mathcal{S}}(Z\mid Y)$ and $P_{\mathcal{T}}(Z\mid Y)$. However, using pseudo labels from model predictions to determine the target class-conditional distributions exposes the alignment to noisy supervision—especially early in training.
Under the Optimal Transport (OT) framework, it has been investigated in some theoretical works that the generalization error on the target domain is controlled by both the marginal alignment loss and the entanglement between the source and target domains. For example, Redko et al. (2017) proves the following:
Theorem 1 (Informal Redko et al. (2017)).
@@
A bad pulling strategy on target domain $\mathcal{T}$ might minimize $W_{1}$ term to 0 without any guarantee for the $\lambda$ term in the feature space. Similarly, Koç et al. (2025) also show that, during the optimal transport association process, the source inputs $x$ can be associated to target inputs $x^{\prime}$ that have different matching labels. Minimizing the marginal Wasserstein distance between such entangled pairs can cause the entanglement term to increase. To address these challenges, we will focus on the following question in this work:
Question: What is the condition on the domain shift so that the target distribution can be aligned back to the source while preserving the correct class labels? Additionally, with only potentially noisy target pseudo-labels available, is there a theoretically guaranteed and computable metric to quantify the degree of violation of this condition?
Formally, we seek conditions under which the OT between the marginals is label-preserving—i.e., it decomposes into per-class OT between the class-conditional marginals. We formalize and prove these conditions in Section 3. Guided by our theoretical analysis under the semi-discrete OT framework (Section 3.2), we propose the OT score—a confidence metric designed to quantify uncertainty in pseudo-labeled target samples. It measures the degree to which the assigned pseudo-label would violate marginal alignment, thereby serving as a diagnostic of class-conditional alignment. As illustrated in Figure 1, the OT score reflects the flexibility of decision boundaries induced by semi-discrete OT alignment, which enables effective uncertainty estimation in the target domain. This allows the algorithm to abstain from classifying samples with high uncertainty. Compared to fully continuous or fully discrete OT formulations, semi-discrete OT is computationally more efficient, especially in high-dimensional spaces and large-scale datasets. A detailed comparison with existing confidence scores is provided in Appendix A.
We also propose two applications of OT score. First, within SFUDA it acts as a training-time reweighting signal: less confident pseudo labels are down-weighted, suppressing harmful updates and improving accuracy. Second, it provides a reliable label-free proxy for target performance: the mean OT score serves as a surrogate for target error, enabling model selection without target labels.
@@
  • We demonstrate two practical uses of the OT score: (i) as a training-time reweighting signal for SFUDA that down-weights less confident target pseudo-labels to suppress harmful updates and improve accuracy; and (ii) as a label-free proxy for target performance, which enables model selection without target labels.
Notation. Given any probability measure $\mu$ and a measurable map $T$ between measurable spaces, $T:\mathcal{X}\longrightarrow\mathcal{Y}$, we denote $T_{\#}\mu$ the pushforward measure on $\mathcal{Y}$ which is characterized by $\left(\mathrm{T}_{\#}\mu\right)(A)=\mu\left(\mathrm{T}^{-1}(A)\right)$ for measurable set $A$. Let $\hat{\mu}$ denote the corresponding empirical measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ where $x_{i}$ are i.i.d. samples from $\mu$. We also write $x\in\hat{\mu}$ to indicate $x\in\{x_{i}\}_{i=1}^{N}$. If not otherwise specified, $\|\cdot\|$ represents the Euclidean norm.
@@ 2.1 Domain Adaptation
Let $\Omega\subseteq\mathbb{R}^{d}$ be the sample space and $\mathcal{P}(\Omega)$ be the set of all probability measures over $\Omega$. In a general supervised learning paradigm for classification problems, we have a labeling function $f_{\bm{\theta}^{*}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ obtained from a parametric family $f_{\bm{\theta}}$ by training on a set of points $\mathbf{X}^{\mathcal{S}}=\{x_{1}^{\mathcal{S}},...,x_{N^{S}}^{\mathcal{S}}\}$ sampled from a source distribution $P_{\mathcal{S}}\in\mathcal{P}(\Omega)$ and corresponding one-hot encoded labels $\mathbf{Y}^{\mathcal{S}}=\{y_{1}^{\mathcal{S}},...,y_{N^{\mathcal{S}}}^{\mathcal{S}}\}$.
Let $\mathbf{X}^{\mathcal{T}}=\{x_{1}^{\mathcal{T}},...,x_{N^{\mathcal{T}}}^{\mathcal{T}}\}$ be a dataset sampled from a target distribution $P_{\mathcal{T}}\in\mathcal{P}(\Omega)$ without label information. The difference between $P_{\mathcal{S}}$ and $P_{\mathcal{T}}$ may lead to a poor performance if we use $f_{\bm{\theta}^{*}}$ for the new classification problem. In order to overcome the challenge of distributional shift, a common way is to decompose a neural network $f_{\bm{\theta}}$ into a feature mapping $\phi_{\bm{v}}$ composed with a classifier $h_{\bm{w}}$ such that $f_{\bm{\theta}}=h_{\bm{w}}\circ\phi_{\bm{v}}$, followed by minimizing the distance between $(\phi_{\bm{v}^{*}})_{\#}P_{\mathcal{S}}$ and $(\phi_{\bm{v}})_{\#}P_{\mathcal{T}}$ so that the target distribution will be aligned with the source distribution in the feature space. Then we may classify target data points based on the optimization result in the feature space. Various choices of divergence objective $D((\phi_{\bm{v}^{*}})_{\#}P_{\mathcal{S}},(\phi_{\bm{v}})_{\#}P_{\mathcal{T}})$ can be utilized. In this work, we focus on the distributional alignment between $(\phi_{\bm{v}^{*}})_{\#}P_{\mathcal{S}}$ and $(\phi_{\bm{v}})_{\#}P_{\mathcal{T}}$ using Wasserstein distance.
@@ 2.2.1 General Theory of OT
$W_{p}(\mu,\nu):=\Bigl(\min_{\gamma\in\Gamma(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma\Bigr)^{\frac{1}{p}},$
@@
$\min_{\gamma\in\Gamma(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma$ (KP)
is referred as the Kantorovitch problem in optimal transport. It is shown by Kantorovich–Rubinstein Duality theorem that (KP) has a dual form (Santambrogio, 2015):
@@
With the dual problem introduced, Brenier (1991) proves Brenier’s theorem, which gives a sufficient condition under which the minimizer of the optimal transport problem is unique and is induced by a map $T=\nabla\phi$ for some convex function $\phi$, i.e. the OT map exists.
Under mild conditions on $\mu$ and $\nu$, Brenier’s theorem is satisfied when $c(x,y)=\|x-y\|_{p}$ for $p>1$. Although there is no guarantee about uniqueness of the optimal transport map when $p=1$, the existence of an optimal transport map can be proved through a secondary variational problem (Santambrogio, 2015):
@@
Let $O(\mu,\nu)$ be the optimal transport plans for the cost $\|x-y\|$ and denote by $K_{p}$ the functional associating to $\gamma\in\mathcal{P}(\Omega\times\Omega)$, the quantity $\int\|x-y\|^{p}\ d\gamma$. Under the usual assumption $\mu\ll\mathcal{L}^{d}$, the secondary variational problem
@@ 2.2.2 Semi-discrete Optimal Transport
$W_{p}^{p}(\mu,\nu)=\max_{\bm{w}\in\mathbb{R}^{m}}\sum_{j=1}^{m}\int_{\mathbb{L}_{\bm{w}}(y_{j})}\bigl(\|x-y_{j}\|^{p}-w_{j}\bigr)\ d\mu+\langle\bm{w},\bm{b}\rangle.$
The optimization problem above can be solved by (stochastic) gradient ascent methods since the $j$-th entry of gradient for the objective function can be computed via $b_{j}-\int_{\mathbb{L}_{\bm{w}}(y_{j})}d\mu$. Once the optimal vector $\bm{w}$ is computed, the optimal transport map $T_{\mu}^{\nu}$ simply maps $x\in\mathbb{L}_{\bm{w}}(y_{j})$ to $y_{j}$ (Peyré et al., 2019). Also, it can be shown such OT map is unique under mild assumptions (Hartmann & Schuhmacher, 2017; Geiß et al., 2013). In the rest of the paper, for any $x\in\operatorname{supp}\mu$ and $y_{j}\in\operatorname{supp}\nu$, we denote $\tilde{d}_{\bm{w}}(x,y_{j}):=\|x-y_{j}\|^{p}-w_{j}$. Convergence properties of semi-discrete optimal transport have been studied extensively; see, e.g., Genevay et al. (2016) and Peyré et al. (2019) for details.
@@ 3 Theoretical Analysis
In this section, we present theoretical insights into the use of OT for addressing DA problems. Complete proofs of all theoretical results are provided in Appendix C. For clarity and tractability, we focus on binary classification tasks. An extension to multiclass classification follows by a one-vs-all reduction. As discussed in Section 2.1, our interest lies in neural network–based DA. To this end, we adopt assumptions inspired by Neural Collapse (Kothapalli, 2022), a prevalent phenomenon observed in well-trained neural networks. The extent to which the target feature distribution conforms to the Neural Collapse structure depends on the severity of the distributional shift between the source and target domains.
@@
Neural collapse (NC) is a phenomenon observed in well-trained neural networks where the learned features of samples belonging to the same class converge to a single point or form tightly clustered structures in the feature space, while the features of different classes become maximally separated. NC emerges while training modern classification DNNs past zero error to further minimize the loss (Papyan et al., 2020). During NC, the class means of the DNN’s last-layer features form a symmetric structure with maximal separation angle, while the features of each individual sample collapse to their class means. This simple structure of the feature layer not only appears beneficial for generalization but also helps in transfer learning and adversarial robustness. There are three main theoretical frameworks proposed to explain the emergence of NC: "Unconstrained Features Model" (Lu & Steinerberger, 2022; Tirer & Bruna, 2022; Ji et al., 2021), "Local Elasticity" (Zhang et al., 2021) and "Neural (Tangent Kernel) Collapse" (Seleznova et al., 2024).
In the following subsection, we focus on the setting where the class-conditional distributions in both the source and target domains are supported on, or concentrated within, bounded subsets of the feature space. Stronger NC in the source representation yields smaller cluster radii, thereby strengthening our results. Under this assumption, we analyze how data clusters are transported by the OT map.
@@ 3.1 Sufficient Conditions for Correct Classification
3.2 Semi-Discrete Setting